# Enhancing Low-Cost Video Editing with Lightweight Adaptors and Temporal-Aware Inversion

### v1

@@ 4.1 Implementation Details
4.2 Dataset
We used the MSR-VTT dataset, comprising 10,000 video clips across 20 categories with 20 English captions each, for a total of about 29,000 unique words, split into 6,513 training, 497 validation, and 2,990 testing clips. Using the OpenAI ChatGPT API, we modified captions to generate new content, and employed a specialized DataLoader to batch early denoised adjacent frames, converting MP4 videos into WebDataset format for optimized training. During inference, adapter weights were loaded into the StableDiffusionPipeline and integrated into the model at a 50% ratio.
Table 1: Comparison of algorithms with and without temporal awareness was conducted using LPIPS, CLIP, and FID metrics on the MSR-VTT Xu et al., (2016) testing set. Adapter parameters were applied during the 0.9-1.0 time intervals of the denoising process. Since text-to-video generation relies solely on textual input without any original source video, there is no FID-based distance between real and generated frames.
@@ 4.2 Dataset
Text2Video-Zero Khachatryan et al., (2023) 0.319 31.38
@@
4.3 Quantitative Results
Superior Performance Across Algorithms and Settings The results in Table 2 demonstrate the versatility and effectiveness of our adapter with a different base diffusion model(SD-XL) and ControlNet conditioning settings. With Canny Edge conditioning, it improves FID from 343.21 to 338.78 and LPIPS from 0.729 to 0.725, enhancing fidelity and perceptual quality. For Human Pose, it boosts CLIP from 29.49 to 33.56, while reducing FID from 365.91 to 362.83 and LPIPS from 0.763 to 0.746, ensuring better semantic alignment. Under Depth Map conditioning, it lowers FID from 343.33 to 339.10, LPIPS from 0.721 to 0.718, and improves CLIP from 29.89 to 31.67, validating the adapter’s ability to enhance visual quality, semantic consistency, and inter-frame coherence across diverse conditions.
Table 2: An ablation study on pretrained ControlNet conditions (controlnet-canny, openpose, depth-sdxl-1.0) with adapter results (yellow) on MSR-VTT human-edited cases for Human Pose demonstrates compatibility and performance gains, with enabled components in green.
@@ 4.3 Quantitative Results
Figure 10: Visualization results obtained by Prompt learner taking different number of shared token (2, 4, 6, 8, 10) training and inference.
Table 1 highlights the compatibility of our adapter with various T2I-based T2V algorithms, showcasing its ability to consistently enhance performance across diverse methods Geyer et al., 2023a ; Chu et al., (2024); Li et al., (2023); Khachatryan et al., (2023) with and without adapter injection demonstrates notable improvements in perceptual quality and text-image alignment: Text2Video-Zero achieves smoother transitions and better prompt adherence with LPIPS reduced from 0.402 to 0.319 and CLIP increased from 28.82 to 31.38; TokenFlow improves motion and texture consistency with a FID drop from 166.72 to 143.21, alongside LPIPS and CLIP gains; Vid2Vid enhances motion alignment with LPIPS decreasing from 0.303 to 0.253 and CLIP increasing from 31.94 to 32.78; and VidToMe improves perceptual quality and semantic alignment with LPIPS reduced from 0.126 to 0.114 and CLIP rising from 25.00 to 28.12. indicating enhanced perceptual quality and semantic alignment.
4.4 Qualitative Results
Figure 5 highlights the pivotal role of the adapter in enhancing video quality across various scenarios and algorithms. In the "lion roaring" example, videos without the adapter suffer from inconsistent facial features and unstable motion across frames, whereas the adapter ensures fluid transitions and a cohesive portrayal of the roaring action. In the "child biking through water" scenario, videos without the adapter exhibit artifacts and temporal inconsistencies, such as distorted water reflections and unnatural motion, which the adapter effectively resolves, creating seamless biking dynamics and realistic water effects. Similarly, in the "walking dog" case, frame flickering and disjointed body movements are apparent without the adapter but are eliminated with its use, resulting in smooth, natural strides. Finally, in the "man surfing" example, the adapter improves semantic alignment, maintaining the surfer’s balance and interaction with the waves, yielding visually coherent and dynamic transitions. These case studies demonstrate the adapter’s effectiveness in achieving temporal stability, and semantic alignment, ensuring high-quality video outputs. In Figure 6 and Figure  7, we present a comparative analysis of the cross-attention maps generated by the stable diffusion base model and the modified model enhanced by our adapter during the image generation process. Our observations indicate that subtle variations in the attention maps substantiate the impact of the adapter on the model’s performance.
Figure 11: Comparison of single-channel feature heatmaps with and without adapter fine-tuning for scenarios ’Two People in Conversation’ and ’Making Tomato Gumbo in the Kitchen,’ showcasing features from the attention layer in UNet blocks 4-11. Labels f1/f2 denote adjacent frames, while t1/t2 represent timesteps (t1=932, t2=941) from the DDPM.
Figure 12: The left panel shows the average cosine similarity rise between interval framelatents during DDIM (the mean value is generated from the full training set of samples), the X-axis is the time step of DDIM, and the Y-axis is in units of 1, showing the consistency of the early noise with the adapter enabled. The right panel shows the jitter in similarity between adjacent frames during training, with the X-axis as the training epoch (1 epoch per 100 steps) and the Y-axis in units of 1, shows stabilty from 0 (no adapter) across training epochs as the adapter is enabled.
Figure 13: Bar chart illustrating the performance of four algorithms across three metrics for various combinations of (training timesteps, inference timesteps). The values 0.5, 0.8, and 0.9 represent the ranges 0.5t–1.0t, 0.8t–1.0t, and 0.9t–1.0t, respectively, where t denotes 1000 timesteps. The number on the left of the tuple indicates the training timesteps, while the number on the right represents the inference timesteps. Note that Text2Video, as a text-to-video generation algorithm, lacks pre-edited videos and therefore does not have an FID metric. The best (training timesteps, inference timesteps) combination for each algorithm is marked with an asterisk (*) at the top of the corresponding bar.
Figure 14: Visual comparison of a 0.5-1.0t training range for temporal-aware loss with a 0.9-1.0t adapter activation in inference range versus setting both to 0.5t (t = 1000).
4.5 Ablation Study
Ablation of Token-Based Semantic Consistency Module. Our ablation study demonstrates that shared tokens are critical for global semantic alignment, with the w Adapter model (integrating Shared Token and Adapter) significantly outperforming the Base model by ensuring consistent frame-level semantic focus on elements like the "swan" and "water" across timesteps, mitigating attention drift, enhancing fine-grained details like neck movements and water ripples, reducing irrelevant background attention, and facilitating smooth, cohesive transitions that integrate motion with ripple effects, showcasing superior semantic consistency and detailed representation. Removing shared tokens severely degrades semantic alignment, as shown in Figure  8, where TokenFlow’s CLIP score drops from 29.4 to 20.1, FID increases from 135.7 to 387.8, and LPIPS rises from 0.14 to 0.17, disrupting global coherence and local detail quality, as further visualized in Figure  9, with the lion’s shape and textures becoming unrecognizable.
The removal of shared tokens significantly hampers semantic consistency, as evident in examples like "A lion is roaring" and "A man is surfing" in Vid2Vid, where the generated outputs lose alignment with the intended textual prompts. Similarly, for VidToMe, phrases like "A car is running" and "A child is riding" fail to maintain cohesion in the generated content, demonstrating a clear limitation when shared tokens are omitted.
Conversely, enhancing unshared tokens, while beneficial for maintaining original video attributes, poses its challenges. It often results in the model becoming overly dependent on the initial video content, reducing its adaptability to novel text-driven prompts. This limitation is evident in cases like "A lion is walking on the grass" in TokenFlow and "A man is running" in Text2Video, where the outputs struggle to effectively incorporate the textual input, indicating reduced flexibility in generating novel content.
Figures  8 and  9 highlight the critical need for balance, as excessive unshared tokens disrupt global semantic consistency and frame transitions, degrading all metrics, including TokenFlow’s CLIP score dropping to 27.8, FID increasing, and LPIPS worsening, underscoring the importance of integrating a well-balanced mix of shared and unshared tokens for both detail fidelity and global coherence. We also investigate the impact of the number of shared tokens in the Prompt Adapter on inference performance. In the main text, we analyze the resultant changes in metrics, and here we provide the relevant visualizations (refer to Figure 10).
Our findings support the hypothesis that an adequate number of shared tokens enhances model performance across various algorithms. As the number of shared tokens increases, video quality stabilizes, and image details become smoother and richer. However, we observe that an excessive number of shared tokens may lead to the loss of textual information; for instance, a configuration with 18 shared tokens in vid2vid more effectively captures the Lego style compared to a configuration with 18 tokens. Consequently, we select 18 shared tokens as the optimal configuration for our final results.
Ablation of Multi-frame Latent Diffusion Model (Unet Adapter). This ablation study is to analyze the role of the Unet adapter in enhancing temporal consistency during video generation. Figure 11 highlights the critical impact of the Unet adapter on temporal feature consistency, as the feature map heatmaps reveal stark contrasts between setups with and without the adapter. Without the adapter, inconsistencies such as structural disparities and reduced intensity alignment across adjacent frames (f1, f2) and timesteps (t1, t2) disrupt temporal transitions, leading to artifacts in video generation. This is further supported by the cosine similarity analysis in Figure 12, which evaluates the temporal coherence of the same video across different timesteps and adjacent frames during training. The left graph illustrates the average cosine similarity between frame embeddings at different timesteps within the same video, comparing models with and without adapter injection. Without the adapter, the similarity demonstrates slower growth and lower final values, indicating weaker temporal coherence. In contrast, the model with the adapter steadily improves, achieving higher cosine similarity and nearing 1.0, highlighting the adapter’s ability to enhance temporal consistency across timesteps.
The activation time period of the Unet adapter is crucial, as shown in Figure 13, which evaluates the effects of temporal-aware loss training ranges and adapter activation ranges during inference. While a 0.5-1.0 training range combined with a 0.5-1.0 inference range achieves the best metrics across CLIP, FID, and LPIPS, Figure 14 reveals trade-offs: the broader 0.5-1.0 range introduces excessive constraints, resulting in over-smoothed details and visual blurriness. To resolve this, a narrower 0.9-1.0 inference range was adopted while keeping the broader 0.5-1.0 training range, striking a balance between temporal consistency and sharp, clear visual outputs.
Figure 15: Comparison of video generation results with bilateral filtering using different kernel sizes in the Stable Diffusion pipline.
Figure 16: Comparison of step and kernel sizes $N(\cdot)$ within bilateral filtering inversion in the Stable Diffusion pipeline. The observed trends are consistent with those in vid2vid and three other algorithms (not shown due to space limitations).
Ablation of Hierarchical-Aware Temporal-Spatial Consistency Module This ablation study analyzes bilateral filtering in the Channel-Dependent Spatially Consistent Denoising Blocks. Figure 15 highlights the critical role of bilateral filtering in improving video quality. In the van, penguin, and rabbit cases, bilateral filtering produces smoother motion and clearer transitions between frames, eliminating the blurred edges and jitter seen without it. For the penguin, the waddling motion becomes more fluid and seamless, while the rabbit’s fur retains natural texture, with consistent motion transitions between jumps, free from artifacts. These visual improvements are supported by the quantitative data in Table 3, where TokenFlow with bilateral inversion shows a decrease in FID from 135.66 to 122.52, a slight increase in CLIP from 29.17 to 29.35, and a reduction in LPIPS from 0.142 to 0.136, confirming that bilateral filtering enhances perceptual quality by reducing noise and improving image consistency.
We also explored the advantages of our inversion method at smaller step sizes, as shown in Figure 16, particularly in the 3-5 step configuration. At these smaller step sizes, our inversion method significantly improves temporal coherence, resulting in smoother transitions between frames. This is especially crucial in video generation, where smaller step sizes reduce noise and enhance fine details by minimizing artifacts like flickering or blurry transitions that often occur at larger step sizes. As a result, the model demonstrates more accurate frame-to-frame consistency, with enhanced details and smoother motion, underscoring the effectiveness of our inversion method at smaller step sizes.
Table 3: Ablation Study: Effects of Inv on Different Algorithms.
@@ 4.5 Ablation Study
4.6 User Study
@@
Figure 17: User study comparison of video generation algorithms with and without adding our adapter.
5 Conclusion and Future Work
To enhance temporal consistency and visual quality in video editing, we presented a prompt-learning adaptor leveraging pre-trained text-to-image diffusion models. We provide a plug-and-play solution that reduces flickering and improves text-to-image alignment, integrating seamlessly into existing workflows. The adaptor combines temporal-aware fine-tuning with a shared/unshared token mechanism to deliver significant improvements in both temporal consistency and video quality at low training costs. This application is compatible with a variety of video editing systems, providing a scalable and efficient alternative to text-driven video editing tools.
Limitations: Although the proposed adapter framework achieves significant improvements in temporal consistency, visual quality, and semantic alignment across different algorithms, it still has some limitations: its adaptability to complex prompts (e.g., significant changes to primary objects) is limited; improvements are less noticeable in cases without obvious flickering; in complex semantic scenarios, the minimal parameters make precise improvements challenging; while the shared token strategy ensures global consistency, it restricts frame-level flexibility; there are occasional discrepancies between quantitative metrics and visual outcomes; and scalability to complex video structures or high-resolution content remains insufficient.
Future work: Future work will enhance adaptability to complex prompts, improve subtle case performance, address complex semantics through parameter optimization and advanced architectures, refine token strategies for consistency and flexibility, align metrics with visual quality, enhance scalability for complex and high-resolution videos, explore adaptive temporal-aware fine-tuning, test diverse algorithms, and develop a generalized, personalized adapter with optimized speed, performance, and visual output for broader applicability and superior user experience.
Acknowledgment
This work is partially supported by Shenzhen Fundamental Research (General Program)(WDZC20231129163533001) and Hunan Provincial Transportation Science and Technology Project (No.202403).

### v3

@@ 4.1. Implementation Details
4.2. Dataset
We used the MSR-VTT dataset (Xu et al., 2016), which has 10,000 video clips across 20 categories and 20 English captions each, resulting in about 29,000 unique words. The dataset is split into 6,513 training, 497 validation, and 2,990 test clips. We modified captions using the OpenAI ChatGPT API to produce additional variations and employed a specialized DataLoader to batch adjacent frames in early denoising steps. The MP4 videos were converted to WebDataset format for faster training. During inference, adapter weights were loaded into the StableDiffusionPipeline and merged at a 50% ratio.
Table 1. Comparison of algorithms with and without temporal awareness was conducted using LPIPS, CLIP, and FID metrics on the MSR-VTT testing set. Adapter parameters were applied during the 0.9-1.0 time intervals of the denoising process. Since text-to-video generation relies solely on textual input without any original source video, there is no FID-based distance between real and generated frames.
@@ 4.2. Dataset
Text2Video-Zero (Khachatryan et al., 2023) 0.319 31.38
@@
4.3. Main Results
Superior Performance Across Algorithms and Settings Table 2 shows that our adapter is effective with different base diffusion models (e.g., SD-XL) and ControlNet conditioning modes. For instance, with Canny Edge conditioning, FID improves from 343.21 to 338.78, and LPIPS from 0.729 to 0.725. With Human Pose conditioning, CLIP scores increase from 29.49 to 33.56, while FID improves from 365.91 to 362.83 and LPIPS from 0.763 to 0.746, indicating enhanced semantic alignment. With Depth Map conditioning, FID decreases from 343.33 to 339.10, LPIPS from 0.721 to 0.718, and CLIP rises from 29.89 to 31.67. These results demonstrate the adapter’s ability to improve visual quality, semantic alignment, and frame-to-frame coherence under diverse conditions.
Table 1 highlights the compatibility of our adapter with various T2I-based T2V algorithms, showcasing its ability to consistently enhance performance across diverse methods (Geyer et al., 2023a; Chu et al., 2024; Li et al., 2023; Khachatryan et al., 2023) with and without adapter injection demonstrates notable improvements in perceptual quality and text-image alignment: Text2Video-Zero achieves smoother transitions and better prompt adherence with LPIPS reduced from 0.402 to 0.319 and CLIP increased from 28.82 to 31.38; TokenFlow improves motion and texture consistency with a FID drop from 166.72 to 143.21, alongside LPIPS and CLIP gains; Vid2Vid enhances motion alignment with LPIPS decreasing from 0.303 to 0.253 and CLIP increasing from 31.94 to 32.78; and VidToMe improves perceptual quality and semantic alignment with LPIPS reduced from 0.126 to 0.114 and CLIP rising from 25.00 to 28.12, indicating enhanced perceptual quality and semantic alignment.
5. Ablation Study
Ablation of Token-Based Semantic Consistency Module. Our ablation study confirms that shared tokens are essential for maintaining global semantic alignment. When the adapter model (including shared tokens) is used, it outperforms the Base model by keeping frame-level semantic focus on key elements such as a “swan” or “water,” thus reducing attention drift. This also enhances details, for instance in neck movements or water ripples, and provides smooth transitions that tie motion and ripple effects together for stronger semantic consistency. Removing shared tokens severely degrades semantic alignment, as shown in Figure 15 (in supplementary material), where TokenFlow’s CLIP score drops from 29.4 to 20.1, FID increases from 135.7 to 387.8, and LPIPS rises from 0.14 to 0.17, disrupting global coherence and local detail quality, as further visualized in Figure 12 (in supplementary material), with the lion’s shape and textures becoming unrecognizable.
Boosting unshared tokens helps preserve attributes of the original video but can limit the model’s flexibility for novel textual prompts. For example, TokenFlow struggles with “A lion is walking on the grass,” and Text2Video with “A man is running,” when unshared tokens are too heavily weighted.
Figures 15 and 12 12 (all in supplementary material) underline that too many unshared tokens degrade both global semantics and frame transitions. TokenFlow’s CLIP score, for instance, declines to 27.8, while FID and LPIPS deteriorate. A suitable mix of shared and unshared tokens is thus necessary for both fine detail and broad semantic integrity. We also vary the number of shared tokens used in the Prompt Adapter; visual results in Figure 13 (in supplementary material) and metric outcomes indicate that 18 shared tokens offer an effective balance. Excessive shared tokens risk losing text-specific information.
The activation time period of the Unet adapter is crucial, as shown in Figure 11 (in supplementary material), which evaluates the effects of temporal-aware loss training ranges and adapter activation ranges during inference. While a 0.5-1.0 training range combined with a 0.5-1.0 inference range achieves the best metrics across CLIP, FID, and LPIPS, Figure 14 (in supplementary material) reveals trade-offs: the broader 0.5-1.0 range introduces excessive constraints, resulting in over-smoothed details and visual blurriness. To resolve this, a narrower 0.9-1.0 inference range was adopted while keeping the broader 0.5-1.0 training range, striking a balance between temporal consistency and sharp, clear visual outputs.
Table 2. An ablation study on pretrained ControlNet conditions (controlnet-canny, openpose, depth-sdxl-1.0) with adapter results (yellow) on MSR-VTT human-edited cases for Human Pose demonstrates compatibility and performance gains, with enabled components in green.
@@ 5. Ablation Study
Table 3. Ablation Study: Effects of Inv on Different Algorithms.
@@
Ablation of Hierarchical-Aware Temporal-Spatial Consistency Module We examine the effect of bilateral filtering within the Channel-Dependent Spatially Consistent Denoising Blocks. Figure 7 (in supplementary material) demonstrates that bilateral filtering reduces blur and jitter in generated frames, improving realism for objects such as penguins and rabbits. In Table 3, TokenFlow’s FID improves from 135.66 to 122.52, CLIP rises from 29.17 to 29.35, and LPIPS drops from 0.142 to 0.136, showing tangible gains in perceptual quality and coherence.
Figure 5. Comparison of step and kernel sizes $N(\cdot)$ within bilateral filtering inversion in the Stable Diffusion pipeline. The observed trends are consistent with those in vid2vid and three other algorithms (not shown due to space limitations).
[…]
We also explored the advantages of our inversion method at smaller step sizes, as shown in Figure 5(in supplementary material), particularly in the 3-5 step configuration. At these smaller step sizes, our inversion method significantly improves temporal coherence, resulting in smoother transitions between frames. This is especially crucial in video generation, where smaller step sizes reduce noise and enhance fine details by minimizing artifacts like flickering or blurry transitions that often occur at larger step sizes. As a result, the model demonstrates more accurate frame-to-frame consistency, with enhanced details and smoother motion, underscoring the effectiveness of our inversion method at smaller step sizes.
5.1. User Study
@@
Figure 6. User study comparison of video generation algorithms with and without adding our adapter.
[…]
More experiment details and the qualitative results are in the supplementary material.
6. Conclusion
We introduced a prompt-learning adapter, GE-Adapter, to improve temporal consistency and visual quality in text-guided video editing using pre-trained text-to-image diffusion models. Our plug-and-play adapter integrates temporal, spatial, and semantic consistency with Baliteral DDIM inversion, reducing flickering and refining text-to-video alignment at minimal training cost. Additionally, it incorporates three key components—Frame-based Temporal Consistency Blocks (FTC Blocks), Channel-dependent Spatial Consistency Blocks (SCD Blocks), and a Token-based Semantic Consistency Module (TSC Module)—to enhance perceptual quality and text-image relevance. This approach is also compatible with diverse video editing systems.
Appendix A Qualitative Results
Figure 9 (in supplementary material) highlights the crucial role of the adapter in enhancing video quality across a range of scenarios and algorithms. In the “lion roaring” example, videos without the adapter show inconsistent facial features and unsteady motion, while the adapter achieves fluid transitions and a coherent depiction of the roaring action. In the “child biking through water” scenario, videos without the adapter suffer from artifacts and temporal inconsistencies, including distorted water reflections and unnatural motion. The adapter addresses these issues effectively, creating smooth biking dynamics and realistic water effects. Similarly, in the “walking dog” example, frame flickering and uneven body movements occur without the adapter, but are eliminated when it is used, producing smoother, more natural strides. Finally, in the “man surfing” illustration, the adapter strengthens semantic alignment by maintaining the surfer’s posture and interaction with the waves, resulting in visually cohesive and dynamic transitions. These case studies demonstrate how the adapter can improve temporal stability and semantic alignment, ensuring high-quality video outputs.
In Figures 10 and 8 (all in supplementary material), we compare cross-attention maps generated by the base Stable Diffusion model and our adapter-enhanced model during the image-generation process. Minor differences in these attention maps underscore the adapter’s influence on the model’s performance.
Figure 7. Comparison of video generation results with bilateral filtering using different kernel sizes in the Stable Diffusion pipline.
[…]
Figure 8. Visualization of attention maps comparing video frames generated with and without the adapter in the Stable Diffusion 1.5 pipeline.
[…]
Figure 9. A comparison of video generation quality across different algorithms, with and without the use of an adapter.
[…]
Figure 10. Visualization of attention maps from Unet’s third upsampling block, comparing base and adapter models across 1000 timesteps. The corresponding timesteps from top to bottom are 1, 541, and 981.
[…]
Figure 11. Bar chart illustrating the performance of four algorithms across three metrics for various combinations of (training timesteps, inference timesteps). The values 0.5, 0.8, and 0.9 represent the ranges 0.5t–1.0t, 0.8t–1.0t, and 0.9t–1.0t, respectively, where t denotes 1000 timesteps. The number on the left of the tuple indicates the training timesteps, while the number on the right represents the inference timesteps. Note that Text2Video, as a text-to-video generation algorithm, lacks pre-edited videos and therefore does not have an FID metric. The best (training timesteps, inference timesteps) combination for each algorithm is marked with an asterisk (*) at the top of the corresponding bar.
[…]
Figure 12. Visualization of video generation models (Vid2Vid, VidToMe, Text2Video, and TokenFlow) using different token strategies: Shared Tokens, Enhanced Unshared Tokens, and Removed Shared Tokens, evaluated across multiple prompts.
[…]
Figure 13. Visualization results obtained by Prompt learner taking different number of shared token (2, 4, 6, 8, 10) training and inference.
[…]
Figure 14. Visual comparison of a 0.5-1.0t training range for temporal-aware loss with a 0.9-1.0t adapter activation in inference range versus setting both to 0.5t (t = 1000).
[…]
Figure 15. Comparison of token configurations—removing shared tokens, using shared tokens, and using enhanced unshared tokens—based on LPIPS, CLIP, and FID metrics.
[…]